OpenStudy (anonymous):

3m - 4(1 - 2m) = 18 , the value of m is ?

OpenStudy (anonymous):

ezzat98 just asked this question.

OpenStudy (anonymous):

m=2

OpenStudy (callisto):

@Beven you need to explain your answer, not just give the answer directly

OpenStudy (callisto):

3m - 4(1 - 2m) = 18 3m -4 + 8m = 18 11m = 18+4 11m = 22 m=2

OpenStudy (anonymous):

3m -4(1-2m)=18 expand the brackets 3m -4 + 8m = 18 +4 to both sides 11m=22 divide both sides by 11 m=22/11 m=2